Numerical Analysis I, Fall 2020 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 10

1. Section 4.2: Problems 5, 10, 12.

2. For section 4.2, problem 12, also show K_{1} = 0 alternatively by assuming the expansion

e = (\frac{2 + h}{2 − h})^{\frac{1}{h}} + C_{1}h^{p_{1}} + ⋯

and find p_{1} numerically using N(h) with h = 0.02 and 0.01.

3. Section 4.3: Next week.

4. Section 4.4: Problems 7, 14, 23, 24, 26.

Problem 7 is a programming assignment. Make good use of the ’for’ or ’do’ loop in matlab/C, or the command ’sum’ in matlab.